# HandProxy: Expanding the Affordances of Speech Interfaces in Immersive Environments with a Virtual Proxy Hand

### v1

@@ Abstract.
Hand interaction, speech interface, interaction proxy, virtual environment
copyright: rightsretainedjournalyear: 2025conference: ; March; 2025isbn: XXX-X-XXXX-XXXX-X/XX/XXccs: Human-centered computing Interactive systems and tools
Figure 1. While XR devices increasingly support dynamic hand interactions, the corresponding speech-based interfaces remain limited (a). While adding speech commands enhance interface capability (b), they lack scalability for diverse hand interactions. We introduce the approach of using a virtual proxy hand that enables users to describe their desired interactions, which the system translates into executable hand movements (c). This approach supports various interactions, such as UI control, object manipulation, and interactions that emerged from physics, collisions, or hand rotation and movement (d).
@@ 1. Introduction
However, users may not always be able to perform the expected hand interactions, making it challenging or even impossible to effectively interact with the virtual environments. This could be caused by situational impairments (Wobbrock, 2019), e.g., when users’ hands are occupied with other physical tasks, they may not be able to manipulate virtual objects and interfaces. Environment constraints could also cause challenges, e.g., users may not be able to move freely in a confined physical space (Kari and Holz, 2023), and virtual objects could be out of physical reach in a enlarged virtual space (Whitlock et al., 2018). Additionally, predefined hand interactions may be inaccessible to users with upper body limitations (Yamagami et al., 2023; Mott et al., 2020). These challenges indicate the need for alternative ways of interacting with virtual environments, specifically those that can reproduce or achieve similar interaction experiences and the expressiveness as the original hand interactions.
@@
Among the various alternatives and enhancements to hand input, speech has emerged as an ideal modality due to its intuitiveness and high degree of freedom. Prior research has explored its applications in object manipulation (Wang et al., 2025), navigation (Hombeck et al., 2023), and environment creation (Aghel Manesh et al., 2024). Additionally, speech has been adopted as an assistive input modality in many mainstream XR devices. For example, Apple Vision Pro provides voice control (app, 2024c) as an alternative to hand input, which can initiate simple interactions, including tap, swipe, and drag and drop on the 2D user interface. Users can also issue high-level commands, such as “turn up the volume”, instead of doing it step-by-step. However, these speech interfaces typically support a limited set of hand interactions and require users to use a rigid, predefined command format to trigger actions. While this approach would work for basic system-level interactions, it falls short in supporting the complex and diverse interactions a hand can do in virtual environments. This limitation is particularly relevant to interactions triggered by hand gestures or movements, such as 3D object manipulation (e.g., pinch, punch, squeeze) and physics-based interactions (e.g., slice a fruit in half with a cutting gesture). Thus, we seek to address the question: “How can we expand the affordances of speech interfaces to support expressive and diverse interactions, especially those that can be achieved via hand interactions in virtual environments?”
@@
Instead of merely designing more speech commands, this work proposes an alternative approach: enabling users to use natural language to control a virtual hand, where it can then simulate corresponding movements and perform necessary interactions on the user’s behalf. This approach is motivated by the observation that many interactions result from a sequence of hand movements, where each step contributes specific meaning to the overall interaction. For example, when pulling a lever, visual effects are triggered as the hand grabs the handle, the corresponding magnitude changes as the hand moves the lever, and the effect finally takes place when the hand releases it. Furthermore, many interactions are not explicitly defined but emerge as a result of other factors, such as collision and physics. For example, an application may not have a specific “clean the table” interaction defined, but the user can achieve this by sweeping the virtual hand across the table, pushing objects away through simulated collision. Therefore, it is impossible to simply create predefined speech commands for all possible interactions, as interactions themselves could be loosely defined, i.e., they are either context-dependent or dynamically generated. These considerations motivate our design choice of preserving the virtual hand within the environment. This allows users to control and reproduce actions as if real hand input is used, enabling them to perform interactions that mirror the capabilities of a physical hand.
Figure 2. Example of how HandProxy would process the user’s speech commands. The user continuously talks to the system, where the system concurrently recognizes the commands, decomposes it into steps, parses it into executable hand control instructions, and calculates the sequence of hand movement to control the virtual hand performing the desired interaction. The processed instructions are added to a queue for continuous hand control. HandProxy utilizes both rule-based methods to detect high-priority commands (e.g., stop, undo) and large language model to interpret natural language input. Whenever a high-priority command is detected (e.g., (5) in the plot), HandProxy will terminate the current execution and perform the high-priority commands.
[…]
@@
To understand how well HandProxy could perform hand manipulations from a diverse range of natural language commands, and to gain insights from users’ experiences of controlling a virtual hand through speech, we conducted a user study with 20 participants. Our study shows that HandProxy enabled users to complete a variety of hand interactions in the virtual environment, including different types of interactions such as mid-air gestures (e.g., swipe left), direct object manipulation (e.g., twist the knob), high-level interaction tasks (e.g., increase the volume), and with varying levels of complexity (e.g., one-step to multi-step interactions). Participants reached a 100% task completion rate, and took an average of 1.09 attempts for their speech commands to be correctly executed by the system (with a command execution accuracy of 91.8%). HandProxy was able to handle diverse variations of participants’ commands for the same tasks, such as using descriptive commands (e.g., “touch your index finger and thumb” to pinch, “grab the red fruit” to grab an apple), varying levels of details (e.g., to increase the volume, either grab the slider and move up, or directly say “maximize the volume”), or sentence structures. Furthermore, participants reported the system to be intuitive, effective, and require minimal learning to use, and pointed out possible improvements including more detailed feedback for enhanced disambiguation and discoverability, greater responsiveness, and supporting additional hand controls.
@@ 2.1. Enhancements and Alternatives for Hand Interactions
Despite these advantages, hand interactions also come with certain limitations. Situational impairments (Wobbrock, 2019) such as occupied hands in AR, physical disabilities (Yamagami et al., 2022, 2023; Carreira et al., 2017) such as upper body limitations, fatigue caused by long-term use of mid-air gestures (Jang et al., 2017; Hincapié-Ramos et al., 2014), or environment constraints such as the confined physical space (Kari and Holz, 2023) or out-of-reach virtual objects (Whitlock et al., 2018) could limit users’ abilities to perform hand input. Additionally, users may have different gesture preferences (Pham et al., 2018; Piumsomboon et al., 2013; Xia et al., 2022), making the standardized gesture input less desirable in certain situations.
@@ 2.2. Speech Interfaces in Immersive Environments
In immersive environments, speech interfaces have been shown to be particularly intuitive, expressive, and natural due to their high degree of freedom (Hepperle et al., 2019; Mcglashan and Axling, 1996; Aslam and Zhao, 2024). As a result, it has emerged as a promising input modality for interactions in virtual environments. Prior work has explored speech-based interactions for a range of tasks, including virtual objects manipulation (Aslam and Zhao, 2024; Bolt, 1980; Sharma et al., 1996; Monteiro et al., 2021), locomotion (Hombeck et al., 2023), and scene creation (Zhang et al., 2024; Aghel Manesh et al., 2024). Commercial XR devices have also integrated speech as a built-in assistive input modality. For example, Apple Vision Pro (app, 2024c), Meta Quest (que, 2024b), and Microsoft HoloLens (hol, 2024b) allow users to issue system commands (e.g., volume adjustment, power control) or perform simple hand gestures (e.g., tap, swipe, drag, and drop) using voice input. Beyond verbal commands, non-verbal speech interactions have also been explored as alternative input methods. These include the use of breathing (Sra et al., 2018), blowing (Zielasko et al., 2015), and sound actions (app, 2024a). While these approaches demonstrate the versatility of speech interfaces, they require users to follow rigid command structures, and the interactions supported are limited to basic hand gestures and system functions.
@@
To address these limitations, our work proposes a flexible speech input method that extends beyond predefined commands. By describing the movement of a proxy virtual hand, we seek to enable expressive interactions in virtual environments, especially those that hands or only hands can do. Additionally, to design an effective speech-based system, we built upon prior work on speech interfaces (Murad et al., 2023; Padmanabha et al., 2024; Hepperle et al., 2019), user expectations for AI-driven agents in VR (Aghel Manesh et al., 2024), and contextual considerations in speech systems (Zargham et al., 2024). These insights inspire the design choices of HandProxy to incorporate embodied knowledge, contextual understanding, conversational memory, interaction state control, and common knowledge integration to enhance the system’s ability to interpret and execute user commands effectively. Combining these design choices, our system can effectively support users to describe their interaction intentions, and the proxy hand can then perform the necessary interactions on the user’s behalf.
@@ 2.3. Interaction Proxies
Interaction proxies are the extra layer inserted between the original and the manifest interface in order to add or modify interactions without changing the app’s source code (Zhang et al., 2017). Their ability to introduce new functionality with minimal modifications makes them particularly valuable for tasks such as input remapping (Zhang et al., 2017; Liang et al., 2023; Zhang et al., 2018; Lu et al., 2024) and UI automation (Li et al., 2017). This concept has also been applied to immersive environments. For example, researchers have created tangible proxies that map physical input with digital interactions (Jain et al., 2023; Hettiarachchi and Wigdor, 2016; Billinghurst et al., 2001; Chulpongsatorn et al., 2023; Dogan et al., 2024). Some work specifically explored the remapping of hand interactions through the proxied interface in virtual environments, such as remapping VR hand interactions to mobile devices (Kari and Holz, 2023) or finger movement (Tseng et al., 2023) to facilitate interactions in constrained spaces. Additionally, McGlashan et al. (Mcglashan and Axling, 1996) introduced proxy agents as an interaction metaphor in VR, allowing users to issue commands to virtual agents that execute tasks on their behalf.
Inspired by these ideas, HandProxy explores how expressive hand interactions can be effectively translated into speech commands. Rather than relying on a rigid set of predefined voice commands for each interaction, HandProxy allows users to control a virtual hand as a proxy, enabling it to perform actions just as a real hand would. This approach ensures that users can issue commands in a natural and intuitive way, while the virtual hand continues to interact with applications as expected. Additionally, this method reduces the need for system-specific modifications to support the mapping and functionality of the speech interface, making this approach compatible and generalizable across different applications that accept hand input.
@@ 3. Primitives of Hand Interaction
Gesture Control: Gesture control defines the fundamental static or dynamic gesture required to initiate a specific interaction, independent of spatial or temporal constraints. For example, to twist an object clockwise, the hand must first assume a grab gesture, then perform an additional rotation to complete the action. These gesture units build the foundation of a hand interaction. In this work, we demonstrate a core set of base gestures commonly used in virtual environment interactions, such as grab, pinch, and cut. For further technical details on extending the gesture set, please refer to section 4.
@@
Given these control primitives—gesture control, target control, spatial control, and temporal control—a hand interaction can be decomposed into one or more combinations of these foundational elements. For example, as shown in Figure 3, the command “pull up” can be broken down into a grab gesture (Gesture Control) and a upward movement (Spatial Control). Similarly, more complex interactions require multiple combinations of these primitives. For instance, the command “put the peach into the basket” involves a grab gesture (Gesture Control), the peach as the interacting target (Target Control), a movement toward the basket (Spatial Control), and a release action (Gesture Control). The command “punch the bag twice” involves a punch gesture (Gesture Control), the bag as the interacting target (Target Control), and a repeat command for 2 times (Temporal Control). Through this structured decomposition, the system can effectively interpret and reproduce a wide array of hand interactions, accommodating both simple commands and more complex interactions. This framework will guide our system design as detailed in section 4, ensuring that our approach remains flexible and intuitive to various interaction needs.
@@ 4. HandProxy: A Speech System for Virtual Environment Interaction Using a Proxy Hand
Figure 4. An overview of the HandProxy system structure. Given the user input, the speech understanding module recognizes the speech into commands, and parses the commands into system-interpretable instructions. The instructions are then sent to hand control module, which generates the corresponding sequence of hand pose to control the movement of the virtual hand. Visualization and feedback module generates appropriate visual feedback for disambiguation and better transparency.
[…]
@@
    Flexible natural speech commands: HandProxy allows users to describe interactions in their own words without rigid formats. It interprets input, decomposes complex commands into executable steps, and prompts for clarification when needed. Users can issue commands at different levels of detail—from precise hand movements (e.g., “pinch” or “grab the box”) to high-level goals (e.g., “increase the brightness” when a brightness knob is present), which brings additional flexibility for user input.
@@
    Real-time streaming input and execution: HandProxy processes speech input in a continuous, streaming fashion, allowing users to speak naturally while the system interprets and executes commands in parallel for a seamless interaction experience. High-priority commands (e.g., “stop”, “undo”) are processed instantly via rule-based methods, while LLMs handle complex instructions. System components run in parallel to minimize delays.
@@
    Feedback for disambiguation and system transparency: To enhance clarity and system transparency, HandProxy provides visual feedback overlays showing recognized commands, expected hand movements, and prompts for retry when needed. When multiple objects match an object description in the command, it shows disambiguation labels to clarify user intent.
As illustrated in Figure 4, HandProxy consists of three major components: speech understanding, hand control, and visualization & feedback. The speech understanding module continuously listens to and processes user input in a streaming fashion in real-time. It segments long speech inputs into individual commands, and adds them to the command queue. The system then interprets each command and decomposes it into a sequence of instructions. Once instructions are available, the hand control module generates a sequence of hand pose data to reproduce the hand movement needed for the interaction. The processed hand data is then sent to the hand state manager, which handles the playback of the hand data and transmits it to the virtual environment to control the virtual hand. Additional visualizations and feedback — including active commands, potential errors, disambiguation hints, and movement path previews — are overlaid onto the user’s viewport, providing additional transparency into the system’s interpretations. The system components run in parallel to increase the efficiency and avoid blocking the hand movement and impacting user experience. These components work together to create a responsive and intuitive interaction system. Next, we discuss the technical details of each system component and conclude with an overview of the design iterations conducted during development, which offers further rationale for our design choices.

### v2

@@ Abstract.
Hand interaction, speech interface, interaction proxy, virtual environment, hand-object interaction
copyright: ccjournal: IMWUTjournalyear: 2025journalvolume: 9journalnumber: 3article: 107publicationmonth: 9doi: 10.1145/3749484conference: Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies; Vol. 9, No. 3, Article 107; September 2025ccs: Human-centered computing Interactive systems and tools
Figure 1. While XR devices increasingly support dynamic hand interactions, the corresponding speech-based interfaces remain limited (a). While adding speech commands enhances interface capability (b), they lack scalability for diverse hand interactions. We introduce the approach of using a virtual proxy hand that enables users to describe their desired interactions, which the system translates into executable hand movements (c). This approach supports various interactions, such as UI control, object manipulation, and interactions that emerge from physics, collisions, or hand rotation and movement (d).
@@
Figure 2. Example of how HandProxy would process the user’s speech commands. The user continuously talks to the system, where the system concurrently recognizes the commands, decomposes them into steps, parses them into executable hand control instructions, and calculates the sequence of hand movements to control the virtual hand performing the desired interaction. The processed instructions are added to a queue for continuous hand control. HandProxy utilizes both rule-based methods to detect high-priority commands (e.g., stop, undo) and large language model to interpret natural language input. Whenever a high-priority command is detected (e.g., (5) in the plot), HandProxy will terminate the current execution and perform the high-priority commands.
[…]
@@ 1. Introduction
However, users may not always be able to perform the expected hand interactions, making it challenging or even impossible to effectively interact with the virtual environments. This could be caused by situational impairments (Wobbrock, 2019), e.g., when users’ hands are occupied with other physical tasks, they may not be able to manipulate virtual objects and interfaces. Environment constraints could also cause challenges, e.g., users may not be able to move freely in a confined physical space (Kari and Holz, 2023), and virtual objects could be out of physical reach in an enlarged virtual space (Whitlock et al., 2018). Additionally, predefined hand interactions may be inaccessible to users with upper body limitations (Yamagami et al., 2023; Mott et al., 2020). These challenges indicate the need for alternative ways of interacting with virtual environments, specifically those that can reproduce or achieve similar interaction experiences and the expressiveness as the original hand interactions.
@@
Among the various alternatives and enhancements to hand input, speech has emerged as an ideal modality due to its intuitiveness and high degree of freedom. Prior research has explored its uses in object manipulation (Wang et al., 2025), navigation (Hombeck et al., 2023), and environment creation (Aghel Manesh et al., 2024). Additionally, speech has been adopted as an assistive input modality in many mainstream XR devices. For example, Apple Vision Pro provides voice control (app, 2024c) as an alternative to hand input, which can initiate simple interactions, including tap, swipe, and drag and drop on 2D interfaces. Users can also issue high-level commands, such as “turn up the volume”, instead of doing it step-by-step. However, these speech interfaces typically support a limited set of hand interactions and require users to use a rigid, predefined command format to trigger actions. While this approach works for basic system-level interactions, it falls short in supporting the complex and diverse interactions a hand can do in virtual environments. This limitation is particularly relevant to interactions triggered by hand gestures or movements, such as 3D object manipulation (e.g., pinch, punch, squeeze) and physics-based interactions (e.g., slice a fruit in half with a cutting gesture). Thus, we seek to address the question: “How can we expand the affordances of speech interfaces to support expressive and diverse interactions, especially those that can be achieved via hand interactions in virtual environments?”
@@
Instead of merely designing more speech commands, this work proposes an alternative approach: enabling users to use natural language to control a virtual hand, where it can then simulate corresponding movements and perform necessary interactions on the user’s behalf. This approach is motivated by the observation that many interactions result from a sequence of hand movements, where each step contributes specific meaning to the overall interaction. For example, when pulling a lever, visual effects are triggered as the hand grabs the handle, the lever moves as the hand moves, and the effect finally takes place when the hand releases it. Furthermore, many interactions are not explicitly defined but emerge as a result of other factors, such as collision and physics. For example, an application may not have a specific “clean the table” interaction defined, but the user can achieve this by sweeping the virtual hand across the table, pushing objects away through simulated collision. Therefore, it is impossible to simply create predefined speech commands for all possible interactions, as interactions themselves could be loosely defined, i.e., they are either context-dependent or dynamically generated. These considerations motivate our design choice of preserving the virtual hand within the environment. This allows users to control and reproduce actions as if real hand input is used, enabling them to perform interactions that mirror the capabilities of a physical hand.
@@
To understand how well HandProxy could perform hand manipulations from diverse natural language commands, and to gain insights from users’ experiences of controlling a virtual hand through speech, we conducted a user study with 20 participants. Our study shows that HandProxy enabled users to complete a variety of hand interactions in the virtual environment, including different types of interactions such as mid-air gestures (e.g., swipe left), direct object manipulation (e.g., twist the knob), high-level interaction tasks (e.g., increase the volume), and with varying levels of complexity (e.g., one-step to multi-step interactions). Participants reached a 100% task completion rate, and took an average of 1.09 attempts for their speech commands to be correctly executed by the system (with a command execution accuracy of 91.8%). HandProxy was able to handle diverse variations of participants’ commands for the same tasks, such as using descriptive commands (e.g., “touch your index finger and thumb” to pinch, “grab the red fruit” to grab an apple), varying levels of details (e.g., to increase the volume, either grab the slider and move up, or directly say “maximize the volume”), or sentence structures. Furthermore, participants reported the system to be intuitive, effective, and require minimal learning to use, and pointed out possible improvements including more detailed feedback for enhanced disambiguation and discoverability, greater responsiveness, and supporting additional hand controls.
@@ 2.1. Enhancements and Alternatives for Hand Interactions
Despite these advantages, hand interactions also come with certain limitations. Situational impairments (Wobbrock, 2019) such as occupied hands in AR, physical disabilities (Yamagami et al., 2022, 2023; Carreira et al., 2017) such as upper body limitations, fatigue caused by long-term use of mid-air gestures (Jang et al., 2017; Hincapié-Ramos et al., 2014), or environment constraints such as the confined physical space (Kari and Holz, 2023) or out-of-reach virtual objects (Whitlock et al., 2018) could limit users’ abilities to perform hand input. Additionally, users may prefer different gesture (Pham et al., 2018; Piumsomboon et al., 2013; Xia et al., 2022), making the standardized gesture input less desirable in certain situations.
@@ 2.2. Speech Interfaces in Immersive Environments
Speech has been widely adopted as a control mechanism across various scenarios. For example, users can issue supported voice commands to interact with mobile devices (Vu et al., 2023; Apple, 2024a) or desktop user interfaces (Igarashi and Hughes, 2001; Apple, 2024b). In robotics, speech is used to control robot actions (House et al., 2009; Telkes et al., 2024). More recently, the integration of speech interfaces with large language models (LLMs) has expanded these capabilities (Vemprala et al., 2023; Padmanabha et al., 2024; Ahn et al., 2022; Fang et al., 2024; Coppari et al., 2025; Zhang et al., 2025; Rifai et al., 2024), enabling functions such as motion planning, task decomposition, and flexible natural language input.
@@
In immersive environments, speech interfaces have been shown to be particularly intuitive, expressive, and natural due to their high degree of freedom (Hepperle et al., 2019; Mcglashan and Axling, 1996; Aslam and Zhao, 2024). As a result, it has emerged as a promising input modality for interactions in virtual environments. Prior work has explored speech-based interactions for various tasks, including virtual objects manipulation (Aslam and Zhao, 2024; Bolt, 1980; Sharma et al., 1996; Monteiro et al., 2021), locomotion (Hombeck et al., 2023), and scene creation (Zhang et al., 2024; Aghel Manesh et al., 2024), using verbal or non-verbal commands such as breathing or sound actions (Sra et al., 2018; Zielasko et al., 2015; app, 2024a). Commercial XR devices have also integrated speech as a built-in assistive input modality. For example, Apple Vision Pro (app, 2024c), Meta Quest (que, 2024b), and Microsoft HoloLens (hol, 2024b) allow users to issue system commands (e.g., volume adjustment, power control) or perform simple hand gestures (e.g., tap, swipe, drag-drop) using speech. While these approaches demonstrate the versatility of speech interfaces, they require users to follow rigid command structures, and the interactions supported are limited to basic hand gestures and system functions.
@@
Recently, integrating large language models (LLMs) into speech interfaces has opened up diverse interaction possibilities, allowing users to directly create, manipulate, query, and engage with their environments through flexible speech input (De La Torre et al., 2024; Wang et al., 2025; Aghel Manesh et al., 2024). Building upon these prior works, we specifically explore ways to expand the affordance of speech interfaces through a generalizable approach, so that it can enable more interaction possibilities, while requiring minimal direct modification of individual applications to support it. To achieve this, we introduce a virtual proxy hand in the pipeline. By describing the movement of a proxy virtual hand in flexible speech input, we seek to enable expressive interactions in virtual environments, especially those that hands or only hands can do. Additionally, to design an effective speech-based system, we built upon prior work on speech interfaces (Murad et al., 2023; Padmanabha et al., 2024; Hepperle et al., 2019), user expectations for AI-driven agents in VR (Aghel Manesh et al., 2024), and contextual considerations in speech systems (Zargham et al., 2024). These insights inspire the design choices of HandProxy to incorporate embodied knowledge, contextual understanding, conversational memory, interaction state control, and common knowledge integration to enhance the system’s ability to interpret and execute user commands effectively. Combining these design choices, our system can effectively support users to describe their interaction intentions, and the proxy hand can then perform the necessary interactions on the user’s behalf.
@@ 2.3. Interaction Proxies
Interaction proxies are the extra layer inserted between the original and the manifest interface to add or modify interactions without changing the app’s source code (Zhang et al., 2017). Their ability to introduce new functionality with minimal modifications makes them particularly valuable for tasks such as input remapping (Zhang et al., 2017; Liang et al., 2023; Zhang et al., 2018; Lu et al., 2024) and UI automation (Li et al., 2017). This concept has also been applied to immersive environments. For example, researchers have created tangible proxies that map physical input with digital interactions (Jain et al., 2023; Hettiarachchi and Wigdor, 2016; Billinghurst et al., 2001; Chulpongsatorn et al., 2023; Dogan et al., 2024). Other work explores remapping of complex 3D input motions in VR to a more accessible ranges of motion or simpler input devices (Wentzel et al., 2025). Some work specifically explored the remapping of hand interactions through the proxied interface in virtual environments, such as remapping VR hand interactions to mobile devices (Kari and Holz, 2023) or finger movement (Tseng et al., 2023) to facilitate interactions in constrained spaces. Additionally, McGlashan et al. (Mcglashan and Axling, 1996) introduced proxy agents as an interaction metaphor in VR, allowing users to issue commands to virtual agents that execute tasks on their behalf.
Inspired by these ideas, HandProxy explores how expressive hand interactions can be effectively translated into speech commands. Rather than relying on a rigid set of predefined voice commands for each interaction, HandProxy allows users to control a virtual hand as a proxy, enabling it to perform actions just as a real hand would. This approach ensures that users can issue commands in a natural and intuitive way, while the virtual hand continues to interact with applications as expected. Also, it reduces the need for system-specific modifications to support the mapping and functionality of the speech interface, making this approach compatible and generalizable across different applications that accept hand input.
@@ 3. Primitives of Hand Interaction
Gesture Control: Gesture control defines the fundamental static or dynamic gesture required to initiate a specific interaction, independent of spatial or temporal constraints. For example, to twist an object clockwise, the hand must first assume a grab gesture, then perform a rotation to complete the action. These gesture units build the foundation of a hand interaction. In this work, we demonstrate a core set of base gestures commonly used in virtual environment interactions, such as grab, pinch, and cut. For further technical details on extending the gesture set, please refer to section 4.
@@
Given these control primitives, a hand interaction can be decomposed into one or more combinations of these foundational elements. As shown in Figure 3, “pull up” can be broken down into a grab gesture (Gesture Control) and a upward movement (Spatial Control). High-level commands that require multiple steps can be decomposed into multiple combinations of these primitives. For instance, “put the apple into the basket” involves three steps – grab the apple, move to the basket, and release it. Through this structured decomposition, the system can effectively interpret and reproduce a wide array of hand interactions, accommodating both simple and more complex interactions. This framework guides our system design as detailed in section 4, ensuring that our approach remains flexible and intuitive to various interaction needs.
@@ 4. HandProxy: A Speech System for Virtual Environment Interaction Using a Proxy Hand
    Flexible natural speech commands: HandProxy allows users to describe interactions in their own words without rigid formats. It interprets input, decomposes complex commands into executable steps, and prompts for clarification if needed. Users can issue commands at different levels of detail—from precise hand movements (e.g., “pinch” or “grab the box”) to high-level goals (e.g., “increase the brightness” when a brightness knob is present), which brings additional flexibility for user input.
@@
    Real-time streaming input and execution: HandProxy takes speech input continuously, allowing users to speak naturally while the system interprets and executes commands in parallel for a seamless interaction experience. High priority commands (e.g., “stop”, “undo”) are processed instantly via rule-based methods, while LLMs handle complex instructions. System components run in parallel to minimize delays.
@@
    Feedback for disambiguation and system transparency: To enhance clarity and transparency, HandProxy provides visual feedback overlays showing recognized commands, expected hand movements, and prompts for retry when needed. When multiple objects match the description in the command, it shows disambiguation labels to clarify user intent.
Below, we begin with an overview of the system design elements, followed by technical details of each components, and conclude with a summary of the design iterations conducted during development to offer rationale for our design choices.
4.1. Design Elements
As illustrated in Figure 4, HandProxy consists of three major components: speech understanding, hand control, and visualization & feedback. The speech understanding module continuously listens to and processes user input in a streaming fashion in real-time using Google Speech-to-text API (Google, 2024b). It segments long speech inputs into individual commands based on punctuation, and adds them to the command queue. Using GPT-4o (OpenAI, 2024), the system then interprets each command and decomposes it into a sequence of instructions in the json format. Once instructions are available, the hand control module uses the hand pose sequences that was either retrieved from FPHAB (Garcia-Hernando et al., 2018) dataset or recorded using LeapMotion and modifies the skeleton data to accommodate movement, rotation, and hand-object interaction. The playback of the hand data is controlled by the hand state manager to support interaction timing control, and is streamed to the target virtual environment for virtual hand control, along with additional visualizations and feedback to be overlaid on top of the user’s view. The system components run in parallel to increase the efficiency and avoid blocking the hand movement and impacting user experience.
The HandProxy system is implemented in Python, with its output streamed to a Unity application via the TCP connection. The Unity application renders a sample 3D virtual environment that contains interactable objects and a virtual hand, simulating an immersive application that supports hand gesture input. It overlays visual feedback specific to HandProxy on top of the immersive environment, as shown in Figure 5. For demonstration and prototyping purposes, the Unity app runs on a laptop, but can be deployed to other supported platforms (e.g., Android on Meta Quest). Demonstrations of the interaction experience on desktop and XR headset setups are provided in the supplementary video.
Figure 4. An overview of the HandProxy system structure. Given the user input, the speech understanding module recognizes the speech into commands, and parses the commands into system-interpretable instructions. The instructions are then sent to the hand control module, which generates the corresponding sequence of hand pose to control the movement of the virtual hand. Visualization and feedback module generates appropriate visual feedback for disambiguation and better transparency.
[…]